Thompson Corporation is planning to issue $130,000, five-year, 6 percent bonds. Interest is payable semi-annually each June 30 and December 31. All of the bonds will be sold on July 1, 2017; they mature on June 30, 2022.

Required:Compute the issue (sale) price on July 1, 2017, if the yield is:  (Round time value factor to 4 decimal places. Round the final answers to the nearest dollar amount.)

a. 6%

b. 5%

c.7%
Answer #1

Solution a:

Chart Values are based on:
n= (5 Years*2) 10 Half years
i= (6%/2) 3.00% Semi annual
Cash Flow Table Value * Amount = Present Value
Principal 0.7441 * $1,30,000 = $96,733
Interest (Annuity) [$130,000*6%*6/12] 8.5302 * $3,900 = $33,268
Price of Bonds $1,30,001

Solution b:

Chart Values are based on:
n= (5 Years*2) 10 Half years
i= (5%/2) 2.50% Semi annual
Cash Flow Table Value * Amount = Present Value
Principal 0.7812 * $1,30,000 = $1,01,556
Interest (Annuity) [$130,000*6%*6/12] 8.7521 * $3,900 = $34,133
Price of Bonds $1,35,689

Solution c:

Chart Values are based on:
n= (5 Years*2) 10 Half years
i= (7%/2) 3.50% Semi annual
Cash Flow Table Value * Amount = Present Value
Principal 0.7089 * $1,30,000 = $92,157
Interest (Annuity) [$130,000*6%*6/12] 8.3166 * $3,900 = $32,435
Price of Bonds $1,24,592